OpenStudy (anonymous):

SQRT80 as whole number and reduced sqrt

jhonyy9 (jhonyy9):

4sqrt5

OpenStudy (saifoo.khan):

factor 80. \[2 \times 2 \times 2 \times 2 \times 5\] \[\sqrt{2 \times 2 \times 2 \times 2 \times 5}\] Take out the pairs, \[2 \times 2 \sqrt 5\]

OpenStudy (saifoo.khan):

therefore, \[4 \sqrt 5\]